OpenStudy (anonymous):

The only solution to x+1=72/x is x = -9

OpenStudy (solomonzelman):

\(\large\color{midnightblue}{ \rm x+1=72/x }\) multiply everything times x. \(\large\color{midnightblue}{ \rm x^2+x=72 }\) \(\large\color{midnightblue}{ \rm x^2+x-72 =0 }\) \(\large\color{midnightblue}{ \rm (x+9)(x-8) =0 }\)

OpenStudy (xguardians):

Correct.

OpenStudy (anonymous):

Take all my medals @SolomonZelman

OpenStudy (xguardians):

TAKE THEM ALL PROFESSOR @SolomonZelman

OpenStudy (anonymous):

thanks :)

OpenStudy (triciaal):

answer is no as shown by solomon 8 and -9